# Leveraging priors on distribution functions for multi-arm bandits

### v1

@@ 1 Introduction
Multi Arm Bandits (MAB) is a paradigmatic framework to study the exploration $\sim$ exploitation dilemma in sequential decision making under uncertainty. Standard algorithms developed within this framework such as Upper-Confidence Bounds (UCB) absed algorithms (Auer et al., 2002) and Thompson sampling (TS) (Thompson, 1933; Russo et al., 2018) have proven to be useful in applications such as clinical trials, ad-placement strategies, etc. However, it remains difficult to apply them to more complicated real world settings such as those arising in agriculture or experimental sciences wherein the underlying uncertainty mechanism is far more sophisticated: the unknown reward distribution corresponding to each arm/action may not even conform to a parametric class of distributions such as the single-parameter exponential family, and usually exhibit characteristics such as multi-modality. With some abuse of terminology, we shall refer to this challenging setting of the MABs as non-parametric MABs, and we report an optimal algorithm for this setting in the current paper.
@@
To begin with, it’s worthwhile to consider the limitations of UCB and Thompson sampling algorithms in some more detail. Firstly, the efficient performance of UCB type algorithms rely on the construction of tight high-probability confidence sequences (Abbasi-Yadkori & Szepesvári, 2011; Auer et al., 2002). However, for complex problems, it becomes difficult to design such sequences, and only approximate confidence sequences can be designed, which generally tend to be statistically suboptimal (Filippi et al., 2010). Next, although Thompson-Sampling (TS)  (Thompson, 1933; Kaufmann et al., 2012) is a neat and elegant Bayesian algorithm, that enjoys the flexibility of incorporating prior knowledge about the bandit environment, it’s efficiency is limited to the regime of conjugate prior/posterior distributions of the relevant scalar/vector parameter, which is generally not possible beyond a few special cases of bandit environments, e.g. Bernoulli, Gaussian. In other regimes, the posterior distributions no longer exhibit a closed form, and require the application of approximate inference schemes such as Markov Chain Monte Carlo (MCMC), variational inference, etc to draw samples from the posterior distributions. This is usually computationally expensive and can easily lead to the suboptimal performance of Thompson sampling (Phan et al., 2019).
@@
In light of the above limitations, one is tempted to look for a statistical-inference technique suitable for handling complicated real-world distribution functions, and finds the answer in Statistical Bootstrap which is a procedure for estimating the distribution of an estimator by resampling (often with replacement) one’s data or a model estimated from the data. Bootstrapping has been widely used as an alternative to statistical-inference based on the assumption of a parametric-model when that assumption is in doubt, or where parametric inference is impossible or requires complicated formulas for the calculation of standard errors.
@@ 3 Background on Dirichlet processes
In the Bayesian formalism (see also section A for more details), an unknown object is treated as a random variable which is then assumed to be drawn from a prior distribution. A Bayesian solution requires developing methods of computation of the posterior distribution from this prior based on available information about the unknown object. When the unknown object is a probability measure (a cumulative distribution function in the present paper, to be precise), one then faces a non-trivial question of how to even define a prior on an infinite dimensional object and also take care of the constraints of a probability measure (sum up to 1 over its support). An elegant solution was offered in Ferguson (1973) wherein the author introduced the idea of a Dirichlet process (DP) – a probability distribution on the space of probability measures which induces finite-dimensional Dirichlet distributions when the data are grouped. To look at it concretely, consider a random probability measure, $G$, on some nice (e.g. Polish) space $\Theta$ (e.g. $\mathbb{R}$). $G$ is said to be DP distributed with base probability measure $F$ (e.g. a Gaussian, Beta, Bernoulli, etc) and concentration parameter $\alpha\in\mathbb{R}^{+}$, denoted as $G\sim\texttt{DP}(\alpha,F)$, if
@@
In the Bayesian parametric framework, one can usually use Bayes rule for deriving posteriors for parametric models, however for non-parametric case, Bayes rule cannot be used in general (see Appendix  A.1 for technical details). Posteriors for some nonparametric priors can be derived utilizing the property of conjugacy. Particularly, an observation model $M\in\mathcal{G}$, and the family of priors $\mathcal{Q}$ are called conjugate if, for any sample size $n$ and any observation sequence $X_{1},...,X_{n}$, the posterior under any prior $Q\in\mathcal{Q}$ is again an element of $\mathcal{Q}$. Also, merely possessing the property of conjugacy is not enough to form a viable Bayesian prior. For example, a generalization of DPs is the so-called Neutral To The Right (NTTR) processes (Dey et al., 2003). Entire family of NTTR is known to be conjugate, but besides the specific case of DPs, there’s no known explicit method of obtaining posterior indices in other members of the NTTR family. This leads us to discuss the form of DP posteriors next.
@@
Let $X_{1}$, … , $X_{n}$ be a sample from an unknown real-valued distribution $G_{0}$ where $X_{i}\in\mathbb{R}$. To estimate $G_{0}$ from a Bayesian perspective (see Appendix A ) we put a prior on the set of all distributions $\mathcal{G}$ and then we compute the posterior distribution of $G_{0}$, given ${\bf X}_{n}=(X_{1},...,X_{n})$. Let’s put a DP prior on the set $\mathcal{G}$. Correspondingly, Let $\texttt{DP}(\alpha,F_{0})$, denote the DP prior. The distribution $F_{0}$ can be thought of as a prior guess at the true distribution $G_{0}$. The number $\alpha$ controls how tightly concentrated the prior is around $F_{0}$. With a DP prior on $G_{0}$, the posterior of $G_{0}$, given ${\bf X}_{n}=(X_{1},...,X_{n})$, enjoys conjugacy, i.e, it is itself a DP given as, $\texttt{DP}(\alpha_{n},\overline{F}_{n})$, where, the posterior indices, $\alpha_{n}$, and $\overline{F}_{n}$ are obtained as follows (Ferguson, 1973; Ghosal, 2010),
@@
where $\delta_{Z_{i}}$ is a discrete measure concentrated at $Z_{i}$, and $q_{i}$ are random weights, generated independent of $Z_{i}$, such that $q_{i}\in[0,1]$, and $\sum_{i=1}^{N}q_{i}=1$. As one can guess, this is analogous to breaking an actual stick into pieces, and hence the name. The author of Sethuraman (1994) reported that if these weights, $q_{i}$, are constructed such that,
@@
and $N$ is $\infty$, then the generated random discrete measure, $P$, in Eq.7 (with $N$ as $\infty$) is such that, $P\sim\texttt{DP}(\alpha,F)$. Ofcourse, for computation one can’t have $N$ as $\infty$, and the infinite series is truncated at some finite $N$, such that a probability mass, $q_{N}=1-\sum_{i=1}^{N-1}q_{i}=\prod_{i=1}^{N}(1-V_{i})$, is put at the last point, $Z_{N}$, and this construction ensures that all weights, $q_{i}$ sum up to one. This finite Stick-breaking representation has been widely used (Ishwaran & James, 2001; Muliere & Tardella, 1998) thanks to its provable optimality in closely approximating the infinite series (see also Appendix B for this and for more details on choosing finite $N$, etc).
@@
With the stick-breaking representation of DP priors at hand, one wonders how to compute DP posteriors in a practically feasible way, and for this, an iterative form of DP posterior comes in handy given as follows (Blackwell & MacQueen, 1973; Sethuraman, 1994),
@@
Here $V_{i}\sim\texttt{Beta}(1,\alpha+i)$, and =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Beginning with a DP prior, $Q_{0}$, generated using the stick-breaking method (Eqs.7-8), the recursion in Eq.9 can be used to obtain the DP posterior, given $N$ observations $\{X_{1},...,X_{N}\}$, as follows,
@@ 4 Dirichlet process posterior sampling
Algorithm 2 Algorithm in Riou & Honda (2020)
1:Horizon $T\geq 1$, number of arms $K\geq 1$
2:for $k=1...K$do
3:     Set $R_{k}:=[1]$, and $N_{k}:=1$
4:end for
5:for $t=1...T$do
6:     for $k=1...K$do
8:     end for
10:     Pull arm $I(t)$ and observe reward $R_{t,I(t)}$.
12:end for
@@ 5 Numerical experiments
In this section, we exhibit empirical performance of DPPS on challenging Bernoulli bandit, Beta bandit, and a real-world agriculture dataset. In the experiments that follow, all regret plots exhibit average regret over 200 independent runs and $10\%-90\%$ quantile levels. For Bernoulli bandits we compare DPPS with Beta-Bernoulli Thompson sampling and UCB. Whereas for the other two environemnts we compare with UCB and a generalized version of Beta/Bernoulli (Agrawal & Goyal, 2013) TS because it’s difficult to implement usual parametric Thompson sampling in those settings (especially for the DSSAT bandit setting). Impressive performance of DPPS in a Gaussian bandit environment (with both mean and variance unknown to the algorithmic agent) is also shown in Sec. C. A discussion on the general choice of (hyper)parameters of DP priors ($\alpha$,$F_{0}$, and truncation level of DP prior) is given in Section D. Corresponding code is provided in the supplementary material.
@@
Computational cost of DPPS
Improved performance and flexibility of DPPS (and other Bootstrap based algorithms such as NPTS) does come with higher computational cost. For example, in the 6-arm Bernoulli bandit environments of horizon $T=10000$, average run-time (over 200 independent runs) of DPPS was around 18 seconds, whereas that of parametric TS (conjugate prior/posterior) was 2-3 seconds. For the 7 arm DSSAT bandit problem, DPPS takes around 20 seconds, NPTS takes around 16 seconds. Sec. E gives a detailed overview of the computational complexity of DPPS. All this said, this run time of DPPS can be significantly brought down by utilizing self-similarity (Ghosal, 2010) of DP posteriors and parallel computation of DP posteriors that a construction exploiting this self-similarity would enjoy, which we plan to do in future.
6 Regret upper bounds for DPPS
In this section, we generalize the information theoretic analysis of Thompson sampling introduced in Russo & Van Roy (2016) to a wider class of probability matching algorithms, and then derive upper bound on Bayesian regret of DPPS. We begin by summarizing the key-steps in the original analysis (Russo & Van Roy, 2016) that are crucial for the aforesaid extension, and also include complete proofs for the sake of completion in Sec. G.
Firstly, the Bayesian regret is re-expressed in terms of the entropy of the posterior distribution of optimal action, and an upper bound on information ratio,
@@ 6 Regret upper bounds for DPPS
For any $T\in N$, provided that $\Gamma_{t}\leq\Gamma$ almost surely for each $t\in{1,..,T}$, $\mathbb{E}\left[\mathrm{Regret}(T,\pi^{TS})\right]\leq\sqrt{\Gamma\mathbb{H}(\alpha_{1})T}$.
The information ratio, $\Gamma_{t}:=\frac{(\mathbb{E}_{t}[R_{t,A_{*}}-R_{t,a}])^{2}}{I_{t}(A^{\star};R_{t,a})}$ is defined as the ratio of the square of the instantaneous expected regret by choosing action $a$ to the instantaneous information gain about optimal action $A^{\star}$ if action $a$ is chosen. Clearly, bounding Bayesian regret of an algorithm boils down to bounding the information-ratio of that algorithm. Particularly, for Thompson-sampling, in $\sigma$-sub-Gaussian reward noise bandit setting, it’s easy to obtain the following bound
Lemma 2.
@@
This bound when combined with Lemma 1 and upper bound of $\log K$ for entropy of any posterior distribution of optimal action leads to the following bound on the Bayesian regret of Thompson sampling,
@@
The proof of Lemma 2 hinges on two crucial steps, and we highlight those, referring the reader to Sec. G for other details. First, re-writing of the instantaneous per-step Bayesian regret by utilizing the probability matching property of Thompson sampling, $\mathbb{P}_{t}(A^{\star}=a)=\mathbb{P}_{t}(A_{t}=a)$, as follows,
@@
If we substitute, the random variable, $X\equiv X(t)=R_{t,a}-\mathbb{E}_{t}[R_{t,a}]$, also the instantaneous reward noise, with $P=\mathbb{P}_{t}(R_{t,a}|A^{\star}=a)$ and $Q=\mathbb{P}_{t}(R_{t,a})$ in the above variational formula, and when $X(t)$ is $\sigma$-sub-Gaussian, it’s easy to obtain the following bound,
@@
6.1 Admissible probability matching algorithms

### v2

@@ 1 Introduction
Multi Arm Bandits (MAB) is a paradigmatic framework to study the exploration $\sim$ exploitation dilemma in sequential decision making under uncertainty. Standard algorithms developed within this framework such as Upper-Confidence Bounds (UCB) based algorithms (Auer et al., 2002) and Thompson sampling (TS) (Thompson, 1933; Russo et al., 2018) have proven to be useful in applications such as clinical trials, ad-placement strategies, etc. However, it remains difficult to apply them to more complicated real world settings such as those arising in agriculture or experimental sciences wherein the underlying uncertainty mechanism is far more sophisticated: the unknown reward distribution corresponding to each arm/action may not even conform to a parametric class of distributions such as the single-parameter exponential family, and usually exhibit characteristics such as multi-modality. With some abuse of terminology, we shall refer to this challenging setting of the MABs as non-parametric MABs, and we report an optimal algorithm for this setting in the current paper.
@@
To begin with, it’s worthwhile to consider the limitations of UCB and Thompson sampling algorithms in some more detail. Firstly, the efficient performance of UCB type algorithms rely on the construction of tight high-probability confidence sequences (Abbasi-Yadkori & Szepesvári, 2011; Auer et al., 2002). However, for complex problems, it becomes difficult to design such sequences, and only approximate confidence sequences can be designed, which generally tend to be statistically suboptimal (Filippi et al., 2010). Next, although Thompson-Sampling (TS)  (Thompson, 1933; Kaufmann et al., 2012) is a neat and elegant Bayesian algorithm, that enjoys the flexibility of incorporating prior knowledge about the bandit environment, it’s efficiency is limited to the regime of conjugate prior/posterior distributions of the relevant scalar/vector parameter, which is generally not possible beyond a few special cases of bandit environments, e.g. Bernoulli, Gaussian. In other regimes, the posterior distributions no longer exhibit a closed form, and require the application of approximate inference schemes such as Markov Chain Monte Carlo (MCMC), variational inference, etc, to draw samples from the posterior distributions. This is usually computationally expensive, and can easily lead to the suboptimal performance of Thompson sampling (Phan et al., 2019).
@@
In light of the above limitations, one is tempted to look for a statistical-inference technique suitable for handling complicated real-world distribution functions, and soon finds the answer in Statistical Bootstrap  (Efron, 1992; Efron & Tibshirani, 1994) which is a procedure for estimating the distribution of an estimator by resampling (often with replacement) data or a model estimated from the data. Bootstrapping has been widely used as an alternative to statistical-inference based on the assumption of a parametric-model when that assumption is in doubt, or where parametric inference is impossible or requires complicated formulas for the calculation of standard errors.
@@ 3 Background on Dirichlet processes
In the Bayesian formalism (see also section 8 for more details), an unknown object is treated as a random variable which is then assumed to be drawn from a prior distribution. A Bayesian solution requires developing methods of computation of the posterior distribution from this prior based on available information about the unknown object. When the unknown object is a probability measure (a cumulative distribution function in the present paper, to be precise), one then faces a non-trivial question of how to even define a prior on an infinite dimensional object and also take care of the constraints of a probability measure (sum up to 1 over its support). An elegant solution was offered in Ferguson (1973) wherein the author introduced the idea of a Dirichlet process (DP) – a probability distribution on the space of probability measures which induces finite-dimensional Dirichlet distributions when the data are grouped. To look at it concretely, consider a random probability measure, $G$, on some nice (e.g. Polish) space $\Theta$ (e.g. $\mathbb{R}$). $G$ is said to be DP distributed with base probability measure $F$ (e.g. a Gaussian, Beta, Bernoulli, etc) and concentration parameter $\alpha\in\mathbb{R}^{+}$, denoted as $G\sim\texttt{DP}(\alpha,F)$, if
@@
In the Bayesian parametric framework, one can usually use Bayes rule for deriving posteriors for parametric models, however for non-parametric case, Bayes rule cannot be used in general (see Appendix  8.1 for technical details). Posteriors for some nonparametric priors can be derived utilizing the property of conjugacy. Particularly, an observation model $M\in\mathcal{G}$, and the family of priors $\mathcal{Q}$ are called conjugate if, for any sample size $n$ and any observation sequence $X_{1},...,X_{n}$, the posterior under any prior $Q\in\mathcal{Q}$ is again an element of $\mathcal{Q}$. Also, merely possessing the property of conjugacy is not enough to form a viable Bayesian prior. For example, a generalization of DPs is the so-called Neutral To The Right (NTTR) processes (Dey et al., 2003). Entire family of NTTR is known to be conjugate, but besides the specific case of DPs, there’s no known explicit method of obtaining posterior indices in other members of the NTTR family. This leads us to discuss the form of DP posteriors next.
@@
Let $X_{1}$, … , $X_{n}$ be a sample from an unknown real-valued distribution $G_{0}$ where $X_{i}\in\mathbb{R}$. To estimate $G_{0}$ from a Bayesian perspective (see Appendix 8 ) we put a prior on the set of all distributions $\mathcal{G}$ and then we compute the posterior distribution of $G_{0}$, given ${\bf X}_{n}=(X_{1},...,X_{n})$. Let’s put a DP prior on the set $\mathcal{G}$. Correspondingly, Let $\texttt{DP}(\alpha,F_{0})$, denote the DP prior. The distribution $F_{0}$ can be thought of as a prior guess at the true distribution $G_{0}$. The number $\alpha$ controls how tightly concentrated the prior is around $F_{0}$. With a DP prior on $G_{0}$, the posterior of $G_{0}$, given ${\bf X}_{n}=(X_{1},...,X_{n})$, enjoys conjugacy, i.e, it is itself a DP given as, $\texttt{DP}(\alpha_{n},\overline{F}_{n})$, where, the posterior indices, $\alpha_{n}$, and $\overline{F}_{n}$ are obtained as follows (Ferguson, 1973; Ghosal, 2010),
@@
where $\delta_{Z_{i}}$ is a discrete measure concentrated at $Z_{i}$, and $q_{i}$ are random weights, generated independent of $Z_{i}$, such that $q_{i}\in[0,1]$, and $\sum_{i=1}^{N}q_{i}=1$. As one can guess, this is analogous to breaking an actual stick into pieces, and hence the name. In a seminal paper, Sethuraman (1994) reported that if these weights, $q_{i}$, are constructed such that,
@@
and $N$ is $\infty$, then the generated random discrete measure, $P$, in Eq.7 (with $N$ as $\infty$) is such that, $P\sim\texttt{DP}(\alpha,F)$. Ofcourse, for computation one can’t have $N$ as $\infty$, and the infinite series is truncated at some finite $N$, such that a probability mass, $q_{N}=1-\sum_{i=1}^{N-1}q_{i}=\prod_{i=1}^{N}(1-V_{i})$, is put at the last point, $Z_{N}$, and this construction ensures that all weights, $q_{i}$ sum up to one. This finite Stick-breaking representation has been widely used (Ishwaran & James, 2001; Muliere & Tardella, 1998) thanks to its provable optimality in closely approximating the infinite series (see also Appendix Section 9 for this and for more details on choosing finite $N$, etc).
@@
With the stick-breaking representation of DP priors at hand, one wonders how to sample from DP posteriors (i.e. $\texttt{DP}(\alpha_{n},\overline{F}_{n})$) in a practically feasible way and, for this, an iterative form of DP posterior (Blackwell & MacQueen, 1973; Sethuraman, 1994) comes in handy, and is given as follows,
@@
Here $V_{i}\sim\texttt{Beta}(1,\alpha+i)$, and =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Beginning with a random measure sampled from the DP prior, $Q_{0}$, generated using the stick-breaking method (Eqs.7-8), the recursion in Eq.9 is simply applied $N$ times (corresponding to $N$ observations $\{X_{1},...,X_{N}\}$) to obtain a random measure, $Q_{N}$, sampled from the DP posterior (i.e. $\texttt{DP}(\alpha_{n},\overline{F}_{n})$),
@@ 5 Numerical experiments
In this section, we exhibit empirical performance of DPPS on challenging Bernoulli bandit, Beta bandit, and a real-world agriculture dataset. In the experiments that follow, all regret plots exhibit average regret over 200 independent runs and $10\%-90\%$ quantile levels. For Bernoulli bandits we compare DPPS with Beta-Bernoulli Thompson sampling and UCB. Whereas for the other two environemnts we compare with UCB and a generalized version of Beta/Bernoulli (Agrawal & Goyal, 2013) TS because it’s difficult to implement usual parametric Thompson sampling in those settings (especially for the DSSAT bandit setting). Impressive performance of DPPS in a Gaussian bandit environment (with both mean and variance unknown to the algorithmic agent) is also shown in Sec. 10. A discussion on the general choice of (hyper)parameters of DP priors ($\alpha$,$F_{0}$, and truncation level of DP prior) is given in Section 11. Corresponding code is provided in the supplementary material, and details on computational costs of DPPS are discussed in Section 12. All simulations were generated using the Python 3 language, and made use of the NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), and Matplotlib (Hunter, 2007) libraries.
@@
6 Information theoretic analysis of DPPS
In Section 4, we saw that, like Thompson sampling, DPPS is a probability matching algorithm. In this section, we utilize this property of DPPS to derive upper bounds on its Bayesian-regret. Particularly, we generalize the information theoretic analysis of Thompson sampling introduced in Russo & Van Roy (2016) to a wider class of probability matching algorithms, and show that DPPS is encapsulated in that generalization. We begin by summarizing the key-steps in the analysis of  Russo & Van Roy (2016) and also include complete proofs for the sake of completion in Sec. 14.
6.1 Information theoretic analysis of Thompson sampling
Russo & Van Roy (2016) introduced an elegant framework for deriving upper bounds on Bayesian regret of Thompson sampling. Firstly, the Bayesian regret is re-expressed in terms of the entropy of the posterior distribution of optimal action, and an upper bound on information ratio,
Theorem 1.
(Russo & Van Roy, 2016) For any $T\in N$, provided that $\Gamma_{t}\leq\Gamma$ almost surely for each $t\in{1,..,T}$, $\mathbb{E}\left[\mathrm{Regret}(T,\pi^{TS})\right]\leq\sqrt{\Gamma\mathbb{H}(\alpha_{1})T}$.
The information ratio, $\Gamma_{t}:=\frac{(\mathbb{E}_{t}[R_{t,A_{*}}-R_{t,a}])^{2}}{I_{t}(A^{\star};R_{t,a})}$, is defined as the ratio of the square of the instantaneous expected regret by choosing action $a$ to the instantaneous information gain about optimal action $A^{\star}$ if action $a$ is chosen. Clearly, bounding Bayesian regret of an algorithm then boils down to bounding the information-ratio of that algorithm. Particularly, for Thompson-sampling, in $\sigma$-sub-Gaussian reward noise bandit setting, one obtains the following bound,
@@ 6.1 Information theoretic analysis of Thompson sampling
(Russo & Van Roy, 2016)
@@
This bound when combined with Theorem 1 and upper bound of $\log K$ for entropy of any posterior distribution of optimal action leads to the following bound on the Bayesian regret of Thompson sampling,
@@
The proof of Lemma 1 hinges on two crucial steps, and we highlight those, referring the reader to Sec. 14 for other details. First, re-writing of the instantaneous per-step Bayesian regret by utilizing the probability matching property of Thompson sampling, $\mathbb{P}_{t}(A^{\star}=a)=\mathbb{P}_{t}(A_{t}=a)$, as follows,
@@
If we substitute, the random variable, $X\equiv X(t)=R_{t,a}-\mathbb{E}_{t}[R_{t,a}]$, i.e. the instantaneous reward noise, with $P=\mathbb{P}_{t}(R_{t,a}|A^{\star}=a)$ and $Q=\mathbb{P}_{t}(R_{t,a})$ in the above variational formula, and when $X(t)$ is $\sigma$-sub-Gaussian, it’s easy to obtain the following bound,
@@
Substituting the result of Lemma 2 in Eq. 11, and utilizing the definition of information gain, $I_{t}(A^{\star};R_{t,a})$, and information ratio, $\Gamma_{t}$, yields the bound in Lemma 1.
6.2 A generalization of the information theoretic analysis in Russo & Van Roy (2016)
Choice of $\mathbb{P}_{t}(A^{\star}=a)$